OpenStudy (anonymous):

hi please help with this question! Need the workings. A copper penny has a mass of 2.6 g Assuming it is pure copper, how much of this mass is contributed by a) the neutrons in the copper nuclei b) the electrons. The answers are a) 1.41 g b) 6.5 x 10^-4

OpenStudy (anonymous):

first you have to calculate no. of moles 2.6 g copper have since 1 mole have Avogadro no. of atoms u can find no. of atoms 2.6 g cu have i think from here you can do it.

OpenStudy (anonymous):

Cu atoms/mole of Cu = 6.040 x 10^23 copper atoms/mole of copper 2.6 g / 63.058 g/mole = 0.04123 moles of Cu 0.04123 mol * 6.040 x 10^23 copper atoms/mole of copper = = 2.490292 * 10^22 copper atoms in 2.6 grams There are 29 electrons per copper atom Mass of 1 electron=9.085*10^(-28) gram Mass of 2.6 gm of Cu electrons =(9.085*10^(-28) g)(29)(2.490292 * 10^22) answer: = 6.561 * 10^-4 ---------------------------

OpenStudy (anonymous):

Copper has a normal atomic mass of 63.546 It's atomic number is 29. It's atomic mass is 63 disregarding other mass in the atom Atomic mass -Atomic number = neutrons mass 29 - 63 = 34 Neutron weight is 34 , a ) answer: 2.6 g of copper * (34 / 63) = 1.4 gram for neutron weight -----------------------------

OpenStudy (anonymous):

I gave the work for answers a) and b) above. Check the work out yourself

OpenStudy (anonymous):

thank you!!